find the possible values for s in the inequality 12s-20<50-3s-25.
Answer by wodkit02(9)

To solve this, pretend at first that the < sign is an = sign.
12s-20 < 50-3s-25 ..... 12s-20 = 50-3s-25
Now collect all s terms on one side of the equation, and get everything else on the opposite side:
12s-20 = 50-3s-25
12s-20+3s = 50-3s+3s-25
15s-20 = 25
15s-20+20 = 25+20
15s = 45
Solving for s, we get s = 3.
Now, put the < back in where the = is to get s < 3.
Let's test this answer:
1 < 3, so put 1 in for s in the ORIGINAL equation
12(1)-20 < 25-3(1)
12-20 < 25-3
-8 < 22
Since this is a true statement, we know that values below 3 work.
Let's test a point greater than 3. If the 12s-20 > 25-3s, then s < 3 solves the original inequality.
4 > 3
12(4)-20 < 25-3(4)
48-20 < 25-12
28 is NOT < 13, so the all values for s must be less than 3.